# Sample and Map from a Single Convex Potential: Generation using Conjugate Moment Measures

### v1

@@ 6 Conclusion
In this work, we borrow inspiration from  (Cordero-Erausquin and Klartag, 2015; Santambrogio, 2016) to define a new representation for probability measures. For any $\rho$, we establish the existence of a convex potential $w$ such that $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. We show that this representation is better suited to generative modelling, since the Gibbs factor $\mathfrak{P}_{w}$ follows more closely the original measure $\rho$, in contrast to Cordero-Erausquin and Klartag’s approach, $\nabla u\,\sharp\,\mathfrak{P}_{u}=\rho$, which results in a Gibbs factor $\mathfrak{P}_{u}$ whose spread is inversely proportional to that of $\rho$. Our conjugate measure factorization provides a single function $w$ to sample noises (using LMC, guaranteed to work with log-concave distributions) and transform these codes using $\nabla w^{*}$. We propose to parameterize the conjugate potential $w$ as an ICNN $w_{\theta}$ and rely on the OT toolbox to define a fixed-point iteration scheme to estimate $w_{\theta}$. While our method demonstrated promising results on generative modeling tasks, its performance is currently hindered by difficulties in training ICNN architecture. Future work may also explore the suitability of replacing $\mathcal{N}(0,I)$ in classic approaches with our pre-trained Gibbs factor $\mathfrak{P}_{w}$.
@@ Acknowledgements
This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-103245).
@@ Appendix A Appendix / Supplementary Material
A.1 Proof of Proposition 1
$T_{m_{A},\Sigma_{A},m_{B},\Sigma_{B}}:x\mapsto m_{B}+M(x-m_{A})$
One get:
By identification, we obtain that $\Delta^{-1}=M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ which leads to $\Delta=\Sigma^{-1}$. It means that the functions in $\{u_{r,\Sigma^{-1}},r\in\mathbb{R}^{d}\}$ are fixed-point of $G_{\rho}$ with $\rho=\mathcal{N}(0,\Sigma)$. According to Cordero-Erausquin and Klartag (2015), the moment potential is unique up to translation, which implies that there are no other convex functions whose moment measure is $\rho$. This observation can be intuitively understood from equation 8 where the term $\ln\left(\int e^{-u}\int e^{-u^{*}}\right)$ is constant, equal to $n\ln(2\pi)$ for centered multivariate normal distributions. Among these distributions, the optimal solution involves selecting $u$ such that $\mathfrak{P}_{u^{*}}=\rho$ in order to minimize the KL divergence term.
A.2 Proof of Proposition 2
The proof is similar to the proof of Proposition 1. Given $\rho=\mathcal{N}(m,\Sigma)$, we are looking for $w$ such that $\nabla w^{*}\#\mathfrak{P}_{w}=\rho$ and $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$. This means that $w(x)=\frac{1}{2}(x-r)^{T}\Delta^{-1}(x-r)$ and $w^{*}(x)=\frac{1}{2}x^{T}\Delta x+x^{T}r$. Moreover $\nabla w^{*}:x\rightarrow\Delta x+r$ must be the OT map between $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$ and $\rho=\mathcal{N}(m,\Sigma)$. The OT map between these two gaussians is known in closed form and is given by $x\rightarrow m+M(x-r)$ with $M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$. Solving $\Delta=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ gives $\Delta=\Sigma^{1/3}$ while solving $r=m-Mr$ gives $r=(I_{d}+\Sigma^{1/3})^{-1}m$. We conclude that $\rho$ is the conjugate moment measure of the unique potential $w$ among those whose Gibbs distribution is Gaussian.
A.3 Proof Theorem 2
@@
A.4 1D Experiments
@@
We conducted two experiments, with $\rho$ defined as a mixture of Gaussian distributions. In the first experiment, $\rho$ is a mixture of four Gaussians: $\mathcal{N}(-0.1,0.07)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.7,0.15)$ with mixture weights $\frac{1}{7},\frac{3}{7},\frac{2}{7},\frac{1}{7}$, respectively. In the second experiment, $\rho$ is a mixture of the three Gaussians $\mathcal{N}(-0.8,0.4)$, $\mathcal{N}(1.5,0.6)$, $\mathcal{N}(3,0.5)$, with weights $\frac{1}{2},\frac{1}{3},\frac{1}{6}$. To approximate these mixtures, we used 400,000 samples and estimated their histograms using 100,000 bins. This allowed us to compute the cumulative distribution and quantile functions using numpy’s cumsum and quantile functions, which were essential for estimating the OT map at each iteration of the fixed-point algorithm. The algorithm converged after 300 iterations. Figure 5 illustrates that, in both experiments, the moment measures $\nabla u\,\sharp\,\mathfrak{P}_{u}$ and the conjugate moment measures $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}$ of the computed potentials perfectly align with $\rho_{i}$ confirming that the fixed-point algorithm successfully identified moment potentials $u$ and conjugate moment potentials $w$ of $\rho_{i}$.
@@
Figure 5: Convergence of the fixed-point algorithm.
A.5 Hyperparameters and infrastructure
@@
For the MNIST experiment, we used MNIST dataset (LeCun et al., 2010) under the terms of the Creative Commons Attribution-Share Alike 3.0 license. In this experiment, we convolve the MNIST images with Gaussian noise of standard deviation $0.3$ to ensure that the OT map from the MNIST distribution to the log-concave distribution $\mathfrak{P}_{w}$ exists, as this is guaranteed when the starting distribution is absolutely continuous (Santambrogio, 2015). This effect is visible in Figure 4.
All runs have been made on single GPU V100 or A100-40GB. All the experiments in Figure 3 were generated using a common set of hyperparameters, detailed in Table 1. The number of particles used (i.e., batch size) and the chosen step size $\gamma$ in the LMC process vary across experiments. For the 2D distributions Circles, S-curve, and Scaled-Rotated S-curve, we used 1024 particles, while for Checkerboard and Diag-Checkerboard, a larger number of particles (4096) was used to ensure the stability of the algorithm. The step size $\gamma$ used in the LMC process depends on the shape of the log concave distribution $\mathfrak{P}_{w_{\theta}^{*}}$. Specifically, we used a step size of 5e-5 for Circles, 1e-4 for S-curve and Scaled-Rotated S-curve, and 1e-3 for Checkerboard and Diag-Checkerboard.

### v2

@@ 6 Conclusion
We borrowed inspiration from  (Cordero-Erausquin and Klartag, 2015) to define a new representation for probability measures. For a given $\rho$, we prove the existence of a convex potential $w$ such that $\nabla w^{*}\,\sharp\,\mathfrak{P}_{w}=\rho$. We show that this representation is better suited to generative modeling, because the Gibbs factor $\mathfrak{P}_{w}$ follows more closely the original measure $\rho$, in contrast to Cordero-Erausquin and Klartag’s approach, $\nabla u\,\sharp\,\mathfrak{P}_{u}=\rho$, which results in a Gibbs factor $\mathfrak{P}_{u}$ whose spread is inversely proportional to that of $\rho$. Our conjugate measure factorization uses $w$ to sample noises (using LMC, guaranteed to work with log-concave distributions) and transforms these codes using $\nabla w^{*}$. We propose to parameterize the conjugate potential $w$ as an ICNN $w_{\theta}$, and estimate it using the OT toolbox in two settings: when the target distribution is accessible via samples, and when it is known up to a normalizing constant. We validate both approaches on generative modeling tasks. In the future, we wish to explore the suitability of replacing $\mathcal{N}(0,I)$ with our pre-trained Gibbs factor $\mathfrak{P}_{w}$ in generative modeling pipelines, and retrain maps on top of it.
@@ Acknowledgements
This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-103245). This work was partially supported by Hi! Paris through the PhD funding of Nina Vesseron.
@@ Appendix A Appendix / Supplementary Material
A.1 Sketch of Proof for Theorem 1
Proof Sketch. The complete proof, along with the theorems used in the proof, can be found in Appendix A.2. We rely on the identity $\nabla w^{*}\circ\nabla w=\mathrm{id}$ that holds for strictly convex functions $w$ and show the existence of a strictly convex potential $w$ that verifies $\nabla w\,\sharp\,\rho=\mathfrak{P}_{w}$. To proceed, we first introduce some notations. Let $\mathcal{L}(\mathbb{R}^{d})$ be the set of functions mapping $\mathbb{R}^{d}$ to {$\mathbb{R}\cup\{+\infty\}$. Given a non-negligible set $\Omega\subset\mathbb{R}^{d}$ and a continuous function $v:\Omega\rightarrow\mathbb{R}$ such that $\int_{\mathbb{R}^{d}}e^{-v(z)}\mathrm{d}z$ is finite, we denote by $\mathfrak{P}_{v}^{\Omega}$ the probability measure with density
$\frac{e^{-v(x)}1_{x\in\Omega}}{\int_{\Omega}e^{-v(z)}\mathrm{d}z}\,.$
We then define, for a probability $\rho$ supported on a non-negligible set $\Omega$, the map $G_{\rho}^{\Omega}:\mathcal{L}(\mathbb{R}^{d})\rightarrow\mathcal{L}(\mathbb{R}^{d})$ which to a potential $v$ associates the Brenier potential, defined in (4), from $\rho$ to $\mathfrak{P}_{v}^{\Omega}$
$G_{\rho}^{\Omega}(v):=\mathfrak{B}(\rho,\mathfrak{P}_{v}^{\Omega})$
Lemma 1.
$\mathcal{M}$ =$=\{f\in\mathscr{C}(\Omega)\;\text{such that}\;\;\forall x,y\in\Omega,\;|f(x)-f(y)|\leq R\|x-y\|_{2}\;\text{and}\;f(0_{\mathbb{R}^{d}})=0\}$
A.2 Proof Lemma 1
@@
A.3 Additional Proof for Proposition 1
$T_{m_{A},\Sigma_{A},m_{B},\Sigma_{B}}:x\mapsto m_{B}+M(x-m_{A})$
One get:
By identification, we obtain that $\Delta^{-1}=M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ which leads to $\Delta=\Sigma^{-1}$. It means that the functions in $\{u_{r,\Sigma^{-1}},r\in\mathbb{R}^{d}\}$ are fixed-point of $G_{\rho}$ with $\rho=\mathcal{N}(0,\Sigma)$. According to Cordero-Erausquin and Klartag (2015), the moment potential is unique up to translation, which implies that there are no other convex functions whose moment measure is $\rho$.
A.4 Proof of Proposition 2
The proof is similar to the additional proof of Proposition 1 provided Section A.3. Given $\rho=\mathcal{N}(m,\Sigma)$, we are looking for $w$ such that $\nabla w^{*}\#\mathfrak{P}_{w}=\rho$ and $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$. This means that $w(x)=\frac{1}{2}(x-r)^{T}\Delta^{-1}(x-r)$ and $w^{*}(x)=\frac{1}{2}x^{T}\Delta x+x^{T}r$. Moreover $\nabla w^{*}:x\rightarrow\Delta x+r$ must be the OT map between $\mathfrak{P}_{w}=\mathcal{N}(r,\Delta)$ and $\rho=\mathcal{N}(m,\Sigma)$. The OT map between these two gaussians is known in closed form and is given by $x\rightarrow m+M(x-r)$ with $M=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$. Solving $\Delta=\Delta^{-1/2}(\Delta^{1/2}\Sigma\Delta^{1/2})^{1/2}\Delta^{-1/2}$ gives $\Delta=\Sigma^{1/3}$ while solving $r=m-Mr$ gives $r=(I_{d}+\Sigma^{1/3})^{-1}m$. We conclude that $\rho$ is the conjugate moment measure of the unique potential $w$ among those whose Gibbs distribution is Gaussian.
Appendix B Experiments in the 1D case where $\rho$ is known from samples
B.1 Proposed algorithm for moment measures
For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$, we define the map $J_{\rho}:\mathcal{L}(\mathbb{R}^{d})\rightarrow\mathcal{L}(\mathbb{R}^{d})$ which to a potential $u$ associates the Brenier potential from $\mathfrak{P}_{u}$ to $\rho$
$J_{\rho}(u):=\mathfrak{B}(\mathfrak{P}_{u},\rho)$
The fixed point of $J_{\rho}$ correspond exactly to the moment potentials of $\rho$. This observation motivates the following fixed-point iteration scheme to compute a moment potential of $\rho$:
$u_{0}:=\tfrac{1}{2}\|\cdot\|^{2};\quad\forall t\geq 1,\quad u_{t+1}:=J_{\rho}(u_{t}),$ (11)
In the 1D case, the OT map between density functions $\mu$ and $\nu$ is known in closed form (see §4.2) which enables the exact application of the fixed-point algorithm.
B.2 Details on the 1D experiments
@@
We conducted two experiments, with $\rho$ defined as a mixture of Gaussian distributions. In the first experiment, $\rho$ is a mixture of four Gaussians: $\mathcal{N}(-0.1,0.07)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.3,0.1)$, $\mathcal{N}(0.7,0.15)$ with mixture weights $\frac{1}{7},\frac{3}{7},\frac{2}{7},\frac{1}{7}$, respectively. In the second experiment, $\rho$ is a mixture of the three Gaussians $\mathcal{N}(-0.8,0.4)$, $\mathcal{N}(1.5,0.6)$, $\mathcal{N}(3,0.5)$, with weights $\frac{1}{2},\frac{1}{3},\frac{1}{6}$. To approximate these mixtures, we used 400,000 samples and estimated their histograms using 100,000 bins. This allowed us to compute the cumulative distribution and quantile functions using numpy’s cumsum and quantile functions, which were essential for estimating the OT map at each iteration of the fixed-point algorithm. The algorithm converged after 300 iterations.
B.3 Code for the 1D experiments
Code for estimating the moment potential in 1D
Listing 1: Code for estimating the moment potential in 1D
@jax.jit
def compute_ot_map(positions_source, freq_source, samples_target):
cdf = jnp.cumsum(freq_source, axis=0)
cdf = cdf/cdf[-1]
quantile_fn = jax.vmap(lambda x: jnp.quantile(samples_target, x))
inverse_cdf = quantile_fn(cdf)
return (positions_source, inverse_cdf, freq_source)
def compute_next_measure(positions_source, freq_source, samples_target):
positions_source, inverse_cdf, freq_source = compute_ot_map(positions_source, freq_source, samples_target)
u = integrate.cumtrapz(inverse_cdf, positions_source, initial=0)
weights = jnp.array(scipy.special.softmax(-u)) * jnp.sum(freq_source)
return (positions_source, weights), (u, inverse_cdf)
# hyperparameters
nb_bins = 100000
bound_min = -11
bound_max = 11
nb_points = 400000
nb_bins_plot = 1000
# generate source and target distributions
rng = jax.random.PRNGKey(0)
rng_source, rng_target, rng = jax.random.split(rng, 3)
mu_target = jnp.array([-0.6, -0.1, 0.3, 0.7]) * 0.8
sigma_target = jnp.array([0.15, 0.07, 0.1, 0.15]) * 0.8
alpha_target = jnp.array((1,3,2,1))
sampler_target = multivariate_gaussians_1D(mu=mu_target, sigma=sigma_target, alpha=alpha_target)
samples_target = sampler_target.generate_samples(rng_target, nb_points)
samples_target = samples_target - jnp.mean(samples_target)
samples_source = jax.random.normal(rng_source, shape=(nb_points,))
# construct bins and histogram
freq_source, edges_source = jnp.histogram(samples_source, nb_bins, range=(bound_min, bound_max))
positions_source = (edges_source[:-1] + edges_source[1:]) / 2.0
freq_target, edges_target = jnp.histogram(samples_target, nb_bins, range=(bound_min, bound_max))
positions_target = (edges_target[:-1] + edges_target[1:]) / 2.0
for i in range(300):
(positions_source, freq_source), (u, grad_u) = compute_next_measure(positions_source, freq_source, samples_target)
Code for estimating the conjugate moment potential in 1D: CMFGen
Listing 2: Code for estimating the conjugate moment potential in 1D
@jax.jit
def compute_ot_map_star(positions_target, freq_target, samples_source):
cdf = jnp.cumsum(freq_target, axis=0)
cdf = cdf/cdf[-1]
quantile_fn = jax.vmap(lambda x: jnp.quantile(samples_source, x))
inverse_cdf = quantile_fn(cdf)
return (positions_target, inverse_cdf, freq_target)
def compute_next_measure(positions_target, freq_target, samples_source):
positions_target, inverse_cdf, freq_target = compute_ot_map_star(positions_target, freq_target, samples_source)
u = integrate.cumtrapz(inverse_cdf, positions_target, initial=0)
weights = jnp.array(scipy.special.softmax(-u)) * jnp.sum(freq_target)
return (positions_target, weights), (u, inverse_cdf)
# hyperparameters
nb_bins = 100000
bound_min = -4
bound_max = 4
nb_points = 400000
nb_bins_plot = 1000
# generate source and target distributions
rng = jax.random.PRNGKey(0)
rng_source, rng_target, rng = jax.random.split(rng, 3)
mu_target = jnp.array([-0.6, -0.1, 0.3, 0.7]) * 0.8
sigma_target = jnp.array([0.15, 0.07, 0.1, 0.15]) * 0.8
alpha_target = jnp.array((1,3,2,1))
sampler_target = multivariate_gaussians_1D(mu=mu_target, sigma=sigma_target, alpha=alpha_target)
samples_target = sampler_target.generate_samples(rng_target, nb_points)
samples_target = samples_target - jnp.mean(samples_target)
samples_source = jax.random.normal(rng_source, shape=(nb_points,))
# construct bins and histogram
freq_source, edges_source = jnp.histogram(samples_source, nb_bins, range=(bound_min, bound_max))
positions_source = (edges_source[:-1] + edges_source[1:]) / 2.0
freq_source_plot, edges_source_plot = jnp.histogram(samples_source, nb_bins)
positions_source_plot = (edges_source_plot[:-1] + edges_source_plot[1:]) / 2.0
freq_target, edges_target = jnp.histogram(samples_target, nb_bins, range=(bound_min, bound_max))
positions_target = (edges_target[:-1] + edges_target[1:]) / 2.0
for i in range(300):
rng_, rng = jax.random.split(rng, 2)
(positions_source, freq_source), (u_star, grad_u_star) = compute_next_measure(positions_target, freq_target, samples_source)
samples_source = jax.random.choice(rng_, positions_source, shape=(nb_points,), p=freq_source)
samples_source = samples_source - jnp.mean(samples_source)
B.4 Other figures
@@
Figure 8: Comparison between the Gibbs factor $\mathfrak{P}_{u}$ and the conjugate Gibbs factor $\mathfrak{P}_{w}$ for two mixtures of 1D Gaussian distributions, $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$. Pushforward densities $\nabla w^{*}\sharp\mathfrak{P}_{w}$ and $\nabla u\sharp\mathfrak{P}_{u}$ closely match the target distributions, illustrating that the fixed-point algorithms converged and succefully estimated the (conjugate) potentials.
Appendix C CMFMA experiments
C.1 1D experiments: Figures
Refer to caption Refer to caption Refer to caption
Figure 9: Energy associated to the mixture of gaussians $\rho_{1}$ (left), learned energy (middle), associated potential (right).
Refer to caption Refer to caption Refer to caption
Figure 10: Energy associated to the mixture of gaussians $\rho_{2}$ (left), learned energy (middle), associated potential (right).
Refer to caption Refer to caption
Figure 11: The pushforward densities $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$ closely match the target distributions $\rho_{1}$ and $\rho_{2}$, illustrating that CMFMA succefully estimated the (conjugate) potentials.
C.2 2D experiments
Figure 12: Learning the conjugate moment potential from an energy. $\mathcal{E}_{1}$, $\mathcal{E}_{2}$ and $\mathcal{E}_{3}$ are learned by regression with CMFSamp. The second column shows the learned energy, the third column displays the corresponding conjugate moment potential, and the fourth column presents samples (in red) drawn from $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$.
The expressions of the three energies used for the 2D experiments with CMFMA are the following:
$\mathcal{E}_{1}(x,y)=-6\cdot\sin\left(\left(\tfrac{3}{10}x\right)^{2}+\left(\tfrac{3}{10}y\right)^{2}\right)$
$\mathcal{E}_{2}(x,y)=\frac{(x^{2}+y-11)^{2}+(x+y^{2}-7)^{2}}{100}-2$
$\mathcal{E}_{3}(x,y)=3\cdot\cos\left(\tfrac{2\pi}{10}x-\tfrac{\pi}{2}\right)\cdot\cos\left(\tfrac{2\pi}{10}y-\tfrac{\pi}{2}\right)$
Appendix D CMFGen experiments
D.1 2D experiments: comparison with generative ICNNs
In this experiment, we compare our method, CMFGen, with the standard ICNN-based generative pipeline (see Amos (2023)) that maps a standard gaussian noise to data. The two approaches are evaluated using the Sinkhorn divergence between generated samples and reference samples from the target distribution, computed over batches of size $N=2048$. We repeat this evaluation across $100$ random seeds applied to both generated and target distributions, and the resulting divergences are summarized by the boxplots in Figure 13. As a baseline, the leftmost column reports the 2-Sinkhorn divergence between two independent samples drawn from the target distribution. The $\varepsilon$ parameter used to compute the Sinkhorn divergence is adapted to the scale of the data. As in Vesseron and Cuturi (2024), we set $\varepsilon=0.05\cdot\mathbb{E}_{x,x^{\prime}\sim\rho}|x-x^{\prime}|^{2}$, where the expectation is estimated from two batches of 2048 samples each, drawn independently from the target distribution $\rho$.
Figure 13: Comparison with generative ICNN across various datasets: Diagonal Checkerboard, Checkerboard, Circles, S-curve, and Scale-Rotated S-curve. Each boxplot shows the 2-Sinkhorn divergences between generated and target samples, computed over $N=2048$ samples and averaged over $100$ random seeds. For each dataset, the first boxplot shows the divergence between two independent samples from the target, the second for the generative ICNN, and the third for our method, CMFGen.
D.2 Cartoon dataset: comparison with generative ICNNs
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14: Additional images for the cartoon experiment using CMFGen. Top left: Images from the cartoon dataset; top right: Transported images using the map $\nabla w$; bottom left: Samples from the Gibbs distribution $\mathfrak{P}_{w_{\theta}}$; bottom right: Samples from the distribution $\nabla w_{\theta}^{*}\sharp\mathfrak{P}_{w_{\theta}}$.
Refer to caption Refer to caption
Figure 15: Images produced with a generative ICNN on the cartoon dataset. Left: Gaussian noise; Right: Generated samples by the ICNN with the same architecture as used by CMFGen, i.e., an ICNN with five hidden layers of size 512 and four quadratic input connections.
D.3 Image Reconstruction: MNIST and Cartoon dataset
Refer to caption              Refer to caption
Figure 16: The learned $w_{\theta}$ trained on the MNIST dataset is used for a post-processing task to recover the masked pixels. Gradient ascent is performed on the masked pixels to maximize the log-probability of the full image. Left: Masked images; Right: Reconstructed images.
Refer to caption Refer to caption
Figure 17: The learned $w_{\theta}$ trained on the Cartoon dataset is used for a post-processing task to recover the masked pixels. Gradient ascent is performed on the masked pixels to maximize the log-probability of the full image. Left: Masked images; Right: Reconstructed images.
Appendix E Hyperparameters and infrastructure
All runs have been made on a single GPU V100-32GB.
Noisy MNIST and Cartoon Datasets
@@ Appendix E Hyperparameters and infrastructure
We convolve the MNIST images with Gaussian noise of standard deviation $0.3$ to ensure the existence of the OT map from the MNIST distribution to the log-concave distribution $\mathfrak{P}_{w}$, as guaranteed when the starting distribution is absolutely continuous (Santambrogio, 2015). Similarly, for the Cartoon dataset, we apply Gaussian noise with a standard deviation of $0.2$ to avoid significantly altering the images.
E.1 Hyperparameters for CMFGen